# Holomorphic Deformations of Hyperbolicity Notions on Compact Complex Manifolds

### v1

@@
Abstract. We investigate deformation properties of balanced hyperbolicity, with particular emphasis on degenerate balanced manifolds and their behavior under modifications.
In this context, we introduce two new notions of hyperbolicity for compact non-Kähler manifolds $X$ of complex dimension $\dim_{\mathbb{C}}X=n$ in degree $1\leq p\leq n-1$, inspired by the work of F. Haggui and S. Marouani on $p$-Kähler hyperbolicity. The first notion, called p-SKT hyperbolicity, generalizes the notions of SKT hyperbolicity and Gauduchon hyperbolicity introduced by S. Marouani. The second notion, called p-HS hyperbolicity, extends the notion of sG hyperbolicity defined by Y. Ma.
We investigate the relationship between these notions of analytic nature and their geometric counterparts, namely Kobayashi hyperbolicity and p-cyclic hyperbolicity for $2\leq p\leq n-1$, and we examine the openness under holomorphic deformations of both $p$-HS hyperbolicity and $p$-Kähler hyperbolicity.
@@ 1 Introduction
In one of his seminal papers [24], M. Gromov introduced a notion of hyperbolicity for compact Kähler manifolds, called Kähler hyperbolicity, defined in terms of a boundedness property relative to a fixed Kähler metric. More precisely, let $(X,\omega)$ be a compact Hermitian manifold with $\dim_{\mathbb{C}}X=n$. A $k$-form $\alpha$ on $X$ is said to be $\widetilde{d}$(bounded) if its pullback $\widetilde{\alpha}:=\pi^{*}\alpha$ on the universal cover $\pi:\widetilde{X}\longrightarrow X$ satisfies $\widetilde{\alpha}=d\beta$ for some $\widetilde{\omega}-$bounded (k1)(k-1)-form $\beta$ on $\widetilde{X}$. The manifold $X$ is called Kähler hyperbolic if it admits a Kähler metric $\omega$ that is $\widetilde{d}$(bounded).
This notion of hyperbolicity has been extended to the non-Kähler setting in several ways. For instance, D. Popovici and S. Marouani introduced balanced hyperbolicity in [38] as a first extension of Gromov’s Kähler hyperbolicity : a compact complex manifold $X$ is said to be balanced hyperbolic if it admits a $\widetilde{d}$(bounded) balanced metric $\omega$ (which includes, for example, Kähler hyperbolic manifolds and another special class called degenerate balanced manifolds). Further generalizations include SKT hyperbolicity and Gauduchon hyperbolicity, defined by S. Marouani in [36], as well as sG hyperbolicity, introduced by Y. Ma in [35].
A first result that we present in this paper is the following :
@@
Let $f:\widehat{X}\longrightarrow X$ be a smooth modification, with $\dim_{\mathbb{C}}X=n\geq 3$. If $\widehat{X}$ is degenerate balanced, then $X$ is again degenerate balanced.
We also study the openness of balanced hyperbolicity and of the class of degenerate balanced manifolds under holomorphic deformations. In this context, we obtain the following result, where the (n1,n)(n-1,n)-th weak $\partial\bar{\partial}$-Lemma property was introduced in [20], Definition 5 :
@@
Let $\{X_{t}:t\in\mathbb{B}_{\varepsilon}(0)\}$ be an analytic family of compact complex $n$-dimensional manifolds such that $X_{0}$ is balanced, and assume that the fibers $X_{t}$ satisfy the (n1,n)(n-1,n)-th weak $\partial\bar{\partial}$-Lemma for $t\neq 0$.
@@
    If $X_{0}$ is degenerate balanced, the fibers $X_{t}$ are also degenerate balanced for small $t$.
One way to generalize these notions of hyperbolicity is by considering conditions in degree $1\leq p\leq n-1$, as done by S. Marouani and F. Haggui in [25], where they introduced the notion of p-Kähler hyperbolicity as a generalization of Kähler hyperbolicity.
In this paper, we introduce the notion of Hermitian-Symplectic (or HS) hyperbolicity, in which the Kähler condition is replaced by the Hermitian-Symplectic condition. We also define degree $p$ analogues of this notion and of SKT hyperbolicity, namely p-SKT hyperbolicity and p-HS hyperbolicity, and we prove the following results :
@@
    If $X$ is $p$-HS hyperbolic or $p$-SKT hyperbolic, then $X$ is $p$-cyclically ((or $p$-dimensional)) hyperbolic ((in the sense of Definition 3.14, [25])).
@@
Theorem D.
Let $\{X_{t}\hskip 1.0pt:\hskip 1.0ptt\in\mathbb{B}_{\varepsilon}(0)\}$ be an analytic family of compact complex $n$-dimensional manifolds, where $\mathbb{B}_{\varepsilon}(0)\subset\mathbb{C}^{N}$ is the open ball of radius $\varepsilon>0$ centred in the origin, and fix a degree $1\leq p\leq n-1$. Then :
@@
2 Deformations of Balanced hyperbolic manifolds
@@
We say that a balanced manifold $X$ is balanced hyperbolic if it admits a d~($\widetilde${d}(bounded)) balanced metric $\omega$, i.e. There exists a bounded (2n3)(2n-3)-form $\alpha$ on $\widetilde{X}$ such that $\widetilde{\omega}^{n-1}=d\alpha$.
@@
    Degenerate balanced manifolds, i.e. $n$-dimensional compact complex manifolds that carry a Hermitian metric $\omega$ such that $\omega^{n-1}$ is $d$-exact (see [46], Proposition 5.4). Obviously, this class of manifolds does not contain any Kähler ones (if $\omega$ is Kähler, $\{\omega^{n-1}\}\neq 0\in H^{2n-2}_{DR}(X,\mathbb{R})$), and examples of such manifolds should appear starting from complex dimension 3. Two categories of examples of degenerate balanced manifolds are known :
@@
    Via smooth modifications of the previous examples. More precisely : Take a smooth modification $f:\widehat{X}\longrightarrow X$, with $\dim_{\mathbb{C}}X=n\geq 3$. Assume that either $\pi_{i}(\widehat{X})=0$ for $2\leq i\leq 2n-3$, or $f$ is a contraction to points, i.e. $f$ restricts to a biholomorphism outside of the preimage of a countable set of points in $X$. Then the following holds :
@@
Proposition 2.2 ([46], Proposition 5.4).
Building on the ideas of [44], Theorem 2.2, we establish the following :
Proposition 2.3.
Let $f:\widehat{X}\longrightarrow X$ be a smooth modification, with $\dim_{\mathbb{C}}X=n\geq 3$. If $\widehat{X}$ is degenerate balanced, then $X$ is again degenerate balanced.
Proof.
Let $E$ be the exceptional divisor of $f$ on $\widehat{X}$, and let $A\subset X$ be the analytic subset such that the restriction $f_{\lvert\widehat{X}\setminus E}:\widehat{X}\setminus E\longrightarrow X\setminus A$ is a biholomorphism. The proof proceeds by contraposition. If $X$ is not degenerate balanced, Proposition 2.2 yields the existence of a non-zero $d$-closed $(1,1)$-current $T\geq 0$. Then $T\underset{\text{loc}}{=}i\partial\bar{\partial}\varphi$ for a locally psh function $\varphi$. In particular, $f^{*}T\underset{\text{loc}}{=}i\partial\bar{\partial}(\varphi\circ f)$, and these local pieces glue together into a globally defined $d$-closed ($1,1$)-current $f^{*}T\geq 0$ on $\widehat{X}$ (see [39] for further details). It remains to check that $f^{*}T$ is not the zero current on $\widehat{X}$ to conclude that $\widehat{X}$ is not degenerate balanced again thanks to Proposition 2.2. Assume, for the sake of contradiction, that $f^{*}T=0$. This would mean that $\text{Supp\hskip 2.0pt}T\subset Y$. We distinguish the following two cases :
Case 1. If all the irreducible components $A_{j}$ of $A$ are such that $\text{codim}_{\mathbb{C}}A_{j}\geq 2$ in $X$, The support theorem [Corollary 2.11 in [13], Chapter III], forces $T=0$, which is a contradiction.
Case 2. If $A$ had certain global irreducible components $A_{j}$ with $\text{codim}_{\mathbb{C}}A_{j}=1$ in $X$, then $T=\underset{j}{\sum}\lambda_{j}[A_{j}]$ with $\lambda_{j}\geq 0$ and $A_{j}]$ is the current of integration on $A_{j}$ thanks to another support theorem [Corollary 2.14 in [13], Chapter III]. So $f^{*}T=0$ implies $\lambda_{j}=0$ $\forall j$, meaning that $T=0$ on $X$, yielding again a contradiction. Hence the desired result follows. ∎
It is natural to ask whether the converse holds as well. Toward answering this question, we first make the following observation :
Observation 2.4.
@@
    $0<\text{Vol}_{\omega}(H)=\dfrac{1}{(n-1)!}\int_{H}\omega^{n-1}=\dfrac{1}{(n-1)!}\int_{H}d\alpha=\dfrac{1}{(n-1)!}\int_{\partial H}\alpha=0$
@@
    Note that the second part of Observation 2.4 does not really require the boundedness of $\alpha$, it only uses the $d$-exactness of $\widetilde{\omega}^{n-1}$.
@@
    We can now infer that the converse of Proposition 2.3 does not hold. Indeed, Let $X$ be a compact degenerate balanced manifold, and consider $\mu:\widehat{X}\longrightarrow X$ to be the blow-up of $X$ at a point $x\in X$. Then the exceptional divisor $E:=\mu^{-1}(x)$ is a complex hypersurface in $\widehat{X}$, hence $\widehat{X}$ is not degenerate balanced due to Observation 2.4.1. Furthermore, $\widehat{X}$ is not even balanced hyperbolic since $E\simeq\mathbb{P}^{n-1}$ (see [38], Proposition 2.10).

### v4

# Deformations of Non-Kähler Hyperbolicity Notions and Modifications of Degenerate Balanced Manifolds

@@
Abstract. We study deformation properties of balanced hyperbolicity, with a particular emphasis on degenerate balanced manifolds and their behavior under smooth modifications.
From a different perspective, we introduce two new notions of hyperbolicity for compact complex non-Kähler manifolds $X$ of complex dimension $\dim_{\mathbb{C}}X=n$, in general degree $2p$ with $1\leq p\leq n-1$. These notions are motivated by the work of D. Popovici and H. Kasuya on partial hyperbolicity in arbitrary degree and by the work of F. Haggui and S. Marouani on $p$-Kähler hyperbolicity. The first notion, called $p$-SKT hyperbolicity, extends SKT hyperbolicity and Gauduchon hyperbolicity to degree $2p$. Similarly, the second notion, called $p$-HS hyperbolicity, generalizes the notion of strongly Gauduchon hyperbolicity introduced by Y. Ma.
We then analyze the relationships between these analytic notions and geometric notions of hyperbolicity, namely Brody/Kobayashi hyperbolicity and $p$-cyclic hyperbolicity in degree $2p$ for $2\leq p\leq n-1$. In addition, we study the behavior of $p$-HS hyperbolicity and $p$-Kähler hyperbolicity under holomorphic deformations, establishing openness results for these properties.
@@ 1 Introduction
In one of his seminal papers [26], M. Gromov introduced a notion of hyperbolicity for compact Kähler manifolds, called Kähler hyperbolicity, defined in terms of a boundedness property relative to Kähler metrics. More precisely, let $(X,\omega)$ be a compact Hermitian manifold with $\dim_{\mathbb{C}}X=n$. A $k$-form $\alpha$ on $X$ is said to be $\widetilde{d}$(bounded) if its pullback $\widetilde{\alpha}:=\pi^{*}\alpha$ to the universal cover $\pi:\widetilde{X}\longrightarrow X$ satisfies $\widetilde{\alpha}=d\beta$ for some $\widetilde{\omega}-$bounded (k1)(k-1)-form $\beta$ on $\widetilde{X}$. This notion clearly does not depend on the choice of the Hermitian metric on $X$, since all Hermtitian metrics on a compact complex manifold are comparable. The manifold $X$ is called Kähler hyperbolic if it admits a Kähler metric $\omega$ that is $\widetilde{d}$(bounded).
Over the last few years, several authors have proposed extensions of this hyperbolicity notion beyond the Kähler setting in several ways. First, D. Popovici and S. Marouani introduced the notion of balanced hyperbolicity in [43] as a generalization of Gromov’s Kähler hyperbolicity: a compact complex manifold $X$ is said to be balanced hyperbolic if it admits a $\widetilde{d}$(bounded) balanced metric $\omega$. This class includes, for example, Kähler hyperbolic manifolds and another special class called degenerate balanced manifolds). Further generalizations in the same spirit include SKT hyperbolicity and Gauduchon hyperbolicity, defined by S. Marouani in [41], as well as sG hyperbolicity, introduced by Y. Ma in [40].
The main goal of this paper is to study the deformation stability of such non-Kähler hyperbolicity notions, and to introduce new hyperbolicity concepts that fit naturally into this framework. But first, we obtain the following result concerning the behavior of degenerate balanced manifolds under modifications:
@@
Let $f:\widehat{X}\longrightarrow X$ be a smooth surjective modification of compact complex manifolds, with $\dim_{\mathbb{C}}X=n\geq 3$. If $\widehat{X}$ is degenerate balanced, then $X$ is degenerate balanced.
We then investigate the openness under holomorphic deformations of balanced hyperbolicity and of the degenerate balanced class of manifolds:
@@
Let $\{X_{t}:t\in\mathbb{B}_{\varepsilon}(0)\}$ be an analytic family of compact complex $n$-dimensional manifolds such that $X_{0}$ is balanced. Assume that the fibers $X_{t}$ satisfy the weak (n1,n)(n-1,n)-$\partial\bar{\partial}$-property for $t\neq 0$.
@@
    If moreover, $X_{0}$ is degenerate balanced and the fibers $X_{t}$ satisy the mild (n,n2)(n,n-2)-$\partial\bar{\partial}$-property for $t\neq 0$, the fibers $X_{t}$ are degenerate balanced for small $t$.
When the family $\{X_{t}:t\in\mathbb{B}_{\varepsilon}(0)\}$ satisfies the $\partial\bar{\partial}$-property, we get the following stronger statement, which is new in the context of degenerate balanced manifolds:
@@
Let $\{X_{t}:t\in\mathbb{B}_{\varepsilon}(0)\}$ be an analytic family of compact $\partial\bar{\partial}$-manifolds of complex dimension $n$ such that $X_{0}$ is balanced.
  1. 1.
    Assume $X_{0}$ is balanced hyperbolic, and let $\omega$ be a balanced hyperbolic metric on $X_{0}$. Then $\omega$ extends to a $\mathscr{C}^{\infty}$-family of balanced metrics $\{\omega_{t}\}_{t\sim 0}$ that are Gauduchon hyperbolic on the respective fibers $X_{t}$ for small $t$.
  2. 2.
    If moreover, $\omega$ is degenerate balanced, the $\mathscr{C}^{\infty}$-family of balanced metrics $\{\omega\}_{t\sim 0}$ on the fibers $\{X_{t}\}_{t\sim 0}$ are degenerate balanced with a $\mathscr{C}^{\infty}$-family of $d$-potentials $\{w_{t}\}_{t\sim 0}$ for small $t$, i.e. $\omega=dw_{0}$ and $\omega_{t}=dw_{t}$, $\forall t\sim 0$.
Other notions of hyperbolicity were introduced by D. Popovici and H. Kasuya in [32], where they initiated the concept of partial hyperbolicity, which constitutes the first hyperbolicity notion formulated in intermediate degree $2p$ with $1\leq p\leq n-1$. In a related direction, a natural way to generalize the previously mentioned metric notions of hyperbolicity is to impose analogous conditions in bidegree ($p,p$). This approach was developed by S. Marouani and F. Haggui in [27], where they introduced the notion of $p$-Kähler hyperbolicity as a degree $p$ generalization of Kähler hyperbolicity.
In this paper, we introduce the notion of Hermitian-symplectic (or HS) hyperbolicity, in which the Kähler condition is replaced by the Hermitian-symplectic condition. This notion is motivated in part by a question raised by J. Streets and G. Tian on the existence of non-Kähler Hermitian–symplectic manifolds ([60], Question 1.7). We investigate the similarities and differences between hyperbolicity properties in the Hermitian–symplectic and Kähler settings. We also define degree $2p$ analogues of this notion and of SKT hyperbolicity, namely p-SKT hyperbolicity and p-HS hyperbolicity, and we prove the following results:
Theorem D.
@@
    If $X$ is $p$-HS hyperbolic or $p$-SKT hyperbolic, then $X$ is $p$-cyclically hyperbolic ((or $p$-hyperbolic in the sense of Definition 2.5, [32])).
@@
Theorem E.
Let $\{X_{t}\hskip 1.0pt:\hskip 1.0ptt\in\mathbb{B}_{\varepsilon}(0)\}$ be an analytic family of compact complex $n$-dimensional manifolds, where $\mathbb{B}_{\varepsilon}(0)\subset\mathbb{C}^{N}$ is the open ball of radius $\varepsilon>0$ centered at the origin, and fix $1\leq p\leq n-1$. Then:
@@
2 Balanced hyperbolic manifolds
@@
We say that a compact complex manifold $X$ is balanced hyperbolic if it admits a balanced metric $\omega$ such that $\omega^{n-1}$ is d~($\widetilde${d}(bounded)), i.e. there exists a $\widetilde{\omega}-$bounded (2n3)(2n-3)-form $\alpha$ on $\widetilde{X}$ such that $\widetilde{\omega}^{n-1}:=\pi^{*}\omega^{n-1}=d\alpha$.
@@
    Degenerate balanced manifolds, i.e. $n$-dimensional compact complex manifolds that carry a Hermitian metric $\omega$ such that $\omega^{n-1}$ is $d$-exact (see [53], Proposition 5.4). Obviously, this class of manifolds does not contain any Kähler ones (if a compact complex manifold $X$ admits a Kähler metric $\omega$, then $\{\omega^{n-1}\}\neq 0\in H^{2n-2}_{DR}(X,\mathbb{R})$). Examples of such manifolds appear starting from complex dimension 3. Two classes of examples of degenerate balanced manifolds are known:
@@
    Via smooth modifications of the previous examples. More precisely: Take a smooth modification $f:\widehat{X}\longrightarrow X$ of compact complex manifolds, with $\dim_{\mathbb{C}}X=n\geq 3$. Assume that either $\pi_{i}(\widehat{X})=0$ for $2\leq i\leq 2n-3$, or $f$ is a contraction to points, i.e. $f$ restricts to a biholomorphism outside of the preimage of a countable set of points in $X$. Then the following holds:
@@
2.1  Modifications of Compact Degenerate Balanced Manifolds
It is natural to ask whether the degenerate balanced property is preserved under smooth modifications. Toward answering this question, we first make the following observation:
Observation 2.2.
@@ 2.1  Modifications of Compact Degenerate Balanced Manifolds
    $\text{Vol}_{\omega}(H)=\dfrac{1}{(n-1)!}\int_{H}\omega^{n-1}=\dfrac{1}{(n-1)!}\int_{H}d\alpha=\dfrac{1}{(n-1)!}\int_{\partial H}\alpha=0\qquad(\text{since }\partial H=\emptyset)$
@@
    Note that Observation 2.2.(2) does not require the boundedness of $\alpha$, it only uses the $d$-exactness of $\widetilde{\omega}^{n-1}$.
@@
    We can now infer that the degenerate balanced property is not stable under smooth modifications. Indeed, let $X$ be a compact degenerate balanced manifold, and consider $\mu:\widehat{X}\longrightarrow X$ to be the blow-up of $X$ at a point $x\in X$. Then the exceptional divisor $E:=\mu^{-1}(x)$ is a complex hypersurface in $\widehat{X}$, hence $\widehat{X}$ is not degenerate balanced due to Observation 2.2.(1). Furthermore, $\widehat{X}$ is not even balanced hyperbolic since $E\simeq\mathbb{P}^{n-1}$ and $\mathbb{P}^{n-1}$ cannot be embedded in $X$. More precisely, we have:
    Proposition 2.4 ([43], Proposition 2.10).
    Let $X$ be a compact balanced hyperbolic manifold. Then there is no holomorphic map $f:\mathbb{P}^{n-1}\longrightarrow X$ such that $f$ is non-degenerate at some point $x\in\mathbb{P}^{1}$.
It is natural to ask about the converse: if $f:\widehat{X}\longrightarrow X$ is a smooth modification of compact complex manifolds with $\widehat{X}$ is degenerate balanced, then is $X$ degenerate balanced ?
Proposition 2.5 ([53], Proposition 5.4).
The first part of Observation 2.2 is actually a direct consequence of Proposition 2.5, since the existence of a complex hypersurface $H\subset X$ naturally gives rise to the current of integration over $H$:
$H]:\mathscr{C}^{\infty}_{n-1,n-1}(X)\longrightarrow\mathbb{C}\qquad,\hskip 23.49976ptu\longmapsto\int_{H}u$
This defines a non-zero $d$-closed positive ($1,1$)-current.
Building on the ideas of [51], Theorem 2.2, we establish the following:
Proposition 2.6.
Let $f:\widehat{X}\longrightarrow X$ be a smooth surjective modification of compact complex manifolds, with $\dim_{\mathbb{C}}X=n\geq 3$. If $\widehat{X}$ is degenerate balanced, then $X$ is again degenerate balanced.
Proof.
Let $E$ be the exceptional divisor of $f$ on $\widehat{X}$, and let $A\subset X$ be the analytic subset such that the restriction $f_{\lvert\widehat{X}\setminus E}:\widehat{X}\setminus E\longrightarrow X\setminus A$ is a biholomorphism. The proof proceeds by contradiction. If $X$ is not degenerate balanced, Proposition 2.5 yields the existence of a non-zero $d$-closed $(1,1)$-current $T\geq 0$. Then there exists an open cover $\{U_{j}\}_{j\in I}$ of $X$ such that $T_{\lvert_{U_{j}}}=i\partial\bar{\partial}\varphi_{j}$ for some plurisubharmonic function $\varphi_{j}:U_{j}\longrightarrow\mathbb{R}\cup\{-\infty\}$. In particular, $f^{*}T_{\lvert_{f^{-1}(u_{j})}}=i\partial\bar{\partial}(\varphi_{j}\circ f)$, and these local pieces glue together into a globally defined $d$-closed ($1,1$)-current $f^{*}T\geq 0$ on $\widehat{X}$ thanks to the surjectivity of $f$ (see [44] for further details). It remains to check that $f^{*}T$ is not the zero current on $\widehat{X}$ to conclude that $\widehat{X}$ is not degenerate balanced again thanks to Proposition 2.5. Assume, for the sake of contradiction, that $f^{*}T=0$. This would mean that $\text{Supp\hskip 2.0pt}T\subset A$. We distinguish the following two cases:
Case 1. If all the irreducible components $A_{j}$ of $A$ are such that $\text{codim}_{\mathbb{C}}A_{j}\geq 2$ in $X$, the support theorem [Corollary 2.11 in [14], Chapter III] forces $T=0$, which is a contradiction.
Case 2. If $A$ had certain global irreducible components $A_{j}$ with $\text{codim}_{\mathbb{C}}A_{j}=1$ in $X$, then $T=\underset{j}{\sum}\lambda_{j}[A_{j}]$ where $\lambda_{j}\geq 0$ and $A_{j}]$ is the current of integration on $A_{j}$ thanks to another support theorem [Corollary 2.14 in [14], Chapter III]. So $f^{*}T=\underset{j}{\sum}\lambda_{j}[f^{-1}(A_{j})]=0$ implies $\lambda_{j}=0$ $\forall j$, meaning that $T=0$ on $X$, yielding again a contradiction. Hence the desired result follows. ∎
2.2  Deformations of Balanced hyperbolic manifolds